# MineLand: Simulating Large-Scale Multi-Agent Interactions with Limited Multimodal Senses and Physical Needs

### v1

@@ Abstract
Conventional multi-agent simulators often assume perfect information and limitless capabilities, hindering the ecological validity of social interactions. We propose a multi-agent Minecraft simulator, MineLand, that bridges this gap by introducing limited multimodal senses and physical needs. Our simulator supports up to 48 agents with limited visual, auditory, and environmental awareness, forcing them to actively communicate and collaborate to fulfill physical needs like food and resources. This fosters dynamic and valid multi-agent interactions. We further introduce an AI agent framework, Alex, inspired by multitasking theory, enabling agents to handle intricate coordination and scheduling. Our experiments demonstrate that the simulator, the corresponding benchmark, and the AI agent framework contribute to more ecological and nuanced collective behavior. The source code of MineLand and Alex is openly available at https://github.com/cocacola-lab/MineLand.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: A panoramic view of one scene in MineLand, consisting of multiple AI agents. Subfigure 3&6 show interactions demonstrating cooperation and competition among several agents. Subfigure 5&2&4 showcases the scenarios where the limited senses, physical needs, and multi-tasking mechanism reflect. In Subfigure 1, an agent is performing a creative task named Exploration. Two agents in the left cave of Subfigure 5 cooperate to finish a programmatic mining task, while agents in Subfigure 3 are carrying out building construction, which is a hybrid task.
@@
Multi-agent simulators have facilitated rich explorations of an interactive artificial society that reflects collective behavior. From sandbox games such as Smallville (Park et al., 2023) to virtual environments (Bates, 1994; Laird & VanLent, 2001; Gong et al., 2023), researchers and practitioners have been building open-world simulators that can carry multi-agent behaviors and navigate complex human relationships for decades. Especially with the advent of Large Language Models (LLMs) and Vision-Language Models (VLMs), numerous multi-agent simulators based on these technologies have been at the forefront in various fields, from fundamental research to practical applications, such as watch-and-help (WAH) task (Zhang et al., 2023), Smallville (Park et al., 2023) and Overcook games (Gong et al., 2023). However, conventional multi-agent simulators are often under the assumption of perfect information and limitless capabilities. These idealized worlds diverge sharply from the messy reality of human interaction. This gap between simulated agents and real-world humans hinders the ecological validity and richness of social interaction within these platforms (Heil, 1983).
@@
We propose MineLand, a multi-agent Minecraft simulator as shown in Figure 1, that bridges this gap by introducing limited multimodal senses and physical needs as primary drivers of agent behavior and interaction. First and foremost, the essence of the MineLand’s features is the ability to handle the maximum number of agents. Compared to two-agent WAH, single-agent MineDojo (Fan et al., 2022) and twenty-five-agent Smallville (Park et al., 2023), our MineLand enables the utilization of forty-eight agents in the mainstream consumer desktop PC (§$\S$5.1). Secondly, our simulator operates under the fundamental assumption (Heil, 1983) that agents possess only limited multimodal senses: partially observable environments, eco-centric perspective, and limited visual and auditory senses. This mirrors real-life social interactions, where visibility and audibility can be affected by factors such as distance, terrain, and context. These limitations restrict information access, forcing agents to actively communicate to compensate for sensory deficiencies. Thirdly, we integrate realistic physical needs into agents. Agents require fundamental physical needs, such as food, sustenance, and resource management, which adds a time-based aspect to their daily routine procedures. This necessitates collaboration and competition for resources, mirroring the complex interplay of cooperation and self-interest observed in human societies (Doyal & Gough, 1984; Alderfer, 1969). By incorporating these three features, our simulator fosters the emergence of dynamic and ecologically valid multi-agent interactions [^1] Ecological validity refers to interactions between agents within a simulated environment that closely resemble real-world human interactions. For example, actions are situated, adaptive, and environmentally constrained..
MineLand supports a wide range of task structures and difficulties, covering harvest, tech tree, combat, survival, creative, construction, and stage performance tasks. Customizing the number of players is supported. We also offer two task modes: cooperative mode and competitive mode, in addition to the default free mode. This makes it an excellent platform for benchmarking emergent multi-agent capabilities. With its flexibility, we design different scenes, scenarios, and settings, including crowd-sourcing 4499 programmatic tasks and 1536 creative tasks (2 times compared to MineDojo), to fully understand the potential of LLM- or VLM-based AI agents. Additionally, we have introduced 18 hybrid tasks that combine the features of programmatic tasks and creative tasks.
@@
To integrate agents into this simulator, we design an AI agent framework - Alex - inspired by Multitasking theory from Cognition field (Salvucci & Taatgen, 2008). Alex allows for simultaneously simulating and executing intricate coordination and scheduling with multiple tasks. With this interactive agent framework, we have obtained the following intriguing findings: (1) Multimodal information enabling agents to perform more appropriate actions (§$\S$5.2); (2) Multitasking mechanism allows for processing multiple tasks by autonomously determining their priority (§$\S$5.3); (3) Limited senses force agents to actively communicate to compensate for sensory deficiencies (§$\S$5.4); (4) Physical needs mirrors real life; agents with physical needs survive longer (§$\S$5.5); (5) Our agents work together more effectively, with a reduced workload per agent (§$\S$5.6).
@@
With contributions in the simulator (§$\S$2), benchmark (§$\S$3) and AI agent (§$\S$4), we push the boundaries of multi-agent simulation by bridging the gap between virtual agents and real-world humans. This not only advances understanding of AI multi-agents but also holds potential for applications in human dynamics, social psychology, robotics, and game design. We anticipate that this work will serve as a useful foundation for the community to create new algorithms and make progress in the field of embodied AI multi-agent systems.
@@ 2 MineLand Simulator
MineLand, inspired by Malmo (Johnson et al., 2016) and MineDojo (Fan et al., 2022), is a Minecraft simulator where players [^2] In this work, we use the term “player” to refer to both human players and AI agents. When we mention “agent”, we specifically mean AI agents. Humans have the option to access the game either through VR or using a keyboard. can explore, and interact with each other as well as the environments. This section dissects the design and implementation of our simulator, focusing on the architecture, observation space, state space, action space and communication.
@@
2.1 Architecture
MineLand enables a large number of players, with a better architectural design and a mechanism to facilitate unlimited large-scale interactions among them. To accomplish this, we suggest an architecture to expand Mineflayer (PrismarineJS, 2023) as shown in Figure 2. Based on this new architecture, MineLand supports up to $48$ agents on a mainstream consumer desktop PC, which is a substantial improvement compared to other Minecraft platforms that can only support up to 2 agents. We conducted relevant experiments in Section 5.1. The architecture consists of three main modules: the bot module (implemented in Python), the environment module (implemented in Java), and the bridge module (implemented in JavaScript).
With these three modules, we construct the base of this simulator and offer many advantages. Most importantly, it is capable of supporting dozens of players entering the game simultaneously. It differs from previous Minecraft simulators (e.g., Malmo, MineRL, MineDojo), all of which necessitate running a Minecraft game client for each player. This client-based approach comes with a notable drawback - it incurs substantial resource costs and most machines cannot handle running large-scale Minecraft game clients concurrently. In contrast, MineLand adopts a different approach. It simplifies each Minecraft client into a single thread, optimizing performance overhead caused by multiple clients. With MineLand, adding one more agent only requires one more thread, allowing support for up to $48$ agents simultaneously accessing Minecraft. Additionally, with the introduction of the Mineflayer, a new mode of execution - code, is offered. In contrast to generating textual action descriptions and training an external controller module (Cai et al., 2023) for transforming the plan to the executable code, this simulator is the first language-model-friendly simulator, providing the code instead of the high-level action description. Through the code, it can directly execute plans generated by the language model. Beyond its language-model-friendly nature, MineLand’s three features, namely large-scale multi-agents, limited multimodal senses, and physical needs, make it ideal for developing multi-agent simulators. agents in MineLand generate a series of codes based on the Mineflayer API while representing planning, such as moving, watching, and mining. The advantage of using code is that it avoids error accumulation, whereas using a textual plan requires an additional model to map the plan to the code, which can lead to error accumulation.
@@ 2.2 Observation Space
Our observation space is designed to be compatible with almost all APIs of the popular MineDojo framework. We provide information related to touch, vision, and sound: tactile information (information about the blocks surrounding the agent, which represent the objects that the agent can touch), auditory information, and visual information (RGB video from the first-person perspective of the agent).
These three modalities together provide the agents with multimodal senses. Note that this information is all raw perceptual information. [^4] Besides the raw perceptual information, MineLand also provides the events encountered by the agent, such as injury, death, and others. Injury events can also be regarded as tactile information, but they are presented in the form of events for simplicity. Additionally, we detail the mechanisms and limitations implemented for vision and hearing, including distance attenuation, environmental obstructions, and directional constraints, to model the limited senses.
@@ 2.3 State Space
Previous simulators focus on task-oriented activities, thus containing state space of inventory and equipment. We design a fascinating state space that blends task-oriented activities with the rhythms of daily life. For the rhythms of daily life, we define the states of agents themselves: physical needs like sleep and hunger. Basic physical needs are the foundation that leads to daily behavior. Here are detailed explanations:
Blending the rhythms of daily life with task-oriented activities is what makes this simulator stand out. Imagine agents waking up in their virtual Minecraft homes, engaging in daily routines like cooking and socializing, but also having defined jobs (e.g., lumberjack, farmer) that involve specific task-oriented activities. This creates a natural flow between daily life and goal-driven behavior, providing a more realistic and nuanced environment for studying agent interactions and complex social dynamics.
@@ 2.4 Action Space
The simulator offers a unique action space encompassing both low-level and high-level actions. For the low-level control, MineLand includes basic actions like walking, running, jumping, and interacting with objects. We also incorporate specific task-related actions like mining resources, chopping wood, or crafting tools. Inspired by Wang et al. (2023), these low-level actions are implemented in the form of code. High-level actions provide encapsulated behaviors that may consist of several or hundreds of low-level actions. High-level actions are suitable for complex tasks that require longer computation times. Imagine agents navigating the world, dodging obstacles, and manipulating tools. These complex tasks generate an action sequence (i.e., code), allowing the simulator to continue executing the action sequence, skip some steps earlier, or be interrupted by some special event.
The execution of the code is divided into several steps, with each step lasting 50-200 milliseconds55550 milliseconds is the minimum time unit in Minecraft. We refer to this minimum time unit as a “tick”.. Before taking a step, the agent is provided with the running states of the previous code, either running, ready, or exceptions. After completing a step, the agent, based on the running states, can choose to either switch to a new action code or continue executing the previous code. This function of choosing is implemented by an automatic gate control system with two gates: New and Resume. New means the agent wants to switch to a new code in the following steps. Resume indicates that the agent wants to continue executing the previous code. In this way, the agent can complete a code that needs to be executed for a long period, or be interrupted at an appropriate time.
@@ 2.5 Communication
Agents in MineLand are allowed to communicate with each other, including auditory information, body language (via visual perception), and sharing information in text media. If an agent wants to communicate with other agents, it can directly chat through Minecraft’s message bar, and MineLand will detect it. Only when the distance between other agents and the sending agent is less than a certain threshold, will other agents receive messages. Most importantly, the new message is allowed to interrupt the executing code and execute this message directly before the previous code has ended. This means that even if an agent is working on a 5-minute extension (such as mining), it is still feasible for other agents to communicate with this working agent. This communication mechanism was not supported in the previous work.
@@ 3 MineLand Benchmark Suite and Dataset
Figure 3: Illustration of Task. We have expanded the number of programmatic tasks and creative tasks by 2 times, compared to MineDojo). Additionally, we have introduced hybrid tasks that combine the features of programmatic tasks and creative tasks. Customizing the number of players is supported. For multi-agents, we provide two modes: cooperative mode and competitive mode.
Benchmark Suite offers a wide and diverse range of tasks. These tasks are divided into three categories: programmatic tasks, creative tasks, and hybrid tasks. For each task, customizing the number of players is supported. Competitive mode can be used to measure the differences in capabilities between different AI agents, as well as to develop adversarial learning algorithms. Next, we will provide a detailed introduction to each type of task. Refer to Appendix B for more details.

### v2

@@ Abstract
While Vision-Language Models (VLMs) hold promise for tasks requiring extensive collaboration, traditional multi-agent simulators have facilitated rich explorations of an interactive artificial society that reflects collective behavior. However, these existing simulators face significant limitations. Firstly, they struggle with handling large numbers of agents due to high resource demands. Secondly, they often assume agents possess perfect information and limitless capabilities, hindering the ecological validity of simulated social interactions. To bridge this gap, we propose a multi-agent Minecraft simulator, MineLand, that bridges this gap by introducing three key features: large-scale scalability, limited multimodal senses, and physical needs. Our simulator supports 64 or more agents. Agents have limited visual, auditory, and environmental awareness, forcing them to actively communicate and collaborate to fulfill physical needs like food and resources. Additionally, we further introduce an AI agent framework, Alex, inspired by multitasking theory, enabling agents to handle intricate coordination and scheduling. Our experiments demonstrate that the simulator, the corresponding benchmark, and the AI agent framework contribute to more ecological and nuanced collective behavior. The source code of MineLand and Alex is openly available at https://github.com/cocacola-lab/MineLand.
Figure 1: A panoramic view of one scene in MineLand, consisting of multiple AI agents. Subfigure 3&6 show interactions demonstrating cooperation and competition among several agents. Subfigure 5&2&4 showcases the scenarios where the limited senses, physical needs, and multitasking mechanism reflect. In Subfigure 1, an agent is performing a creative task named Exploration. Two agents in the left cave of Subfigure 5 cooperate to finish a programmatic mining task, while agents in Subfigure 3 are carrying out building construction, which is a hybrid task.
@@ 1 Introduction
Multi-agent simulators have facilitated rich explorations of an interactive artificial society that reflects collective behavior. From sandbox games such as Smallville [40] to virtual environments [4, 32, 18], researchers and practitioners have been building open-world simulators that can carry multi-agent behaviors and navigate complex human relationships for decades. Especially with the advent of Large Language Models (LLMs) and Vision-Language Models (VLMs), numerous multi-agent simulators based on these technologies have been at the forefront in various fields, from fundamental research to practical applications, such as watch-and-help (WAH) task [65], Smallville [40] and Overcook games [18]. However, conventional multi-agent open-world simulators, valuable for exploring collective behavior, suffer from limitations (detailed comparison in Table 1). Firstly, they struggle with large-scale scenarios due to the enormous resource consumption required for large-scale agents. Secondly, they are often under the assumption of perfect information and limitless capabilities. These idealized worlds diverge sharply from the messy reality of human interaction. This gap between existing simulators and the real world hinders the ecological validity and richness of social interaction [21].
@@
To bridge this gap, we propose MineLand (Section 2), a multi-agent Minecraft simulator, as shown in Figure 1, by introducing three key features: large-scale scalability, limited multimodal senses, and physical needs. First and foremost, the essence of the MineLand’s features is the ability to handle the maximum number of agents. Compared to two-agent WAH, single-agent MineDojo [16] and twenty-five-agent Smallville [40], our MineLand enables the utilization of sixty-four or even more agents on the mainstream consumer desktop PC. Secondly, our simulator operates under the human-like assumption [21] that agents possess only limited multimodal senses: partially observable environments, eco-centric perspective, and limited visual and auditory senses. This mirrors real-life social interactions, where visibility and audibility can be affected by factors such as distance, terrain, and environment. These limitations restrict information access, forcing agents to actively communicate to compensate for sensory deficiencies. Thirdly, we integrate realistic physical needs into agents. Agents require fundamental physical needs, such as food, sustenance, and resource management, which adds a time-based aspect to their daily routine procedures. This necessitates collaboration and competition for resources, mirroring the complex interplay of cooperation and self-interest observed in human societies [15, 1]. By incorporating these three features, our simulator fosters the emergence of dynamic and ecologically valid multi-agent interactions [^1] Ecological validity refers to interactions between agents within a simulated environment that closely resemble real-world human interactions. For example, actions are situated, adaptive, and environmentally constrained..
As an open-world multi-agent simulator, MineLand is an excellent platform for benchmarking multi-agent capabilities (Section 3), hence we crowd-sourced datasets to fully evaluate the potential of LLM- or VLM-based multi-agents. For previously existing tasks, programmatic tasks (e.g., harvest, tech tree, combat tasks) and creative tasks (e.g., survival tasks), we offer a significantly expanded task quantity. Specifically, we crowd-sourced 4499 programmatic tasks and 1536 creative tasks, which is 2 times compared to MineDojo. Interestingly, we introduce a novel “hybrid task” category that combines the features of programmatic tasks and creative tasks. Construction and Stage Performance tasks exemplify this new category, bringing the total to 18 hybrid tasks. Additionally, the simulator allows customization of the number of agents and facilitates exploration through two distinct modes: cooperative and competitive.
@@
In addition, we design an AI agent framework - Alex - inspired by Multitasking theory from the Cognition field [53] (Section 4). Alex allows for simultaneously executing intricate coordination and scheduling with multiple tasks. With this AI agent framework, we have obtained the following intriguing findings: (1) Our simulator shows the feature of large-scale scalability (supporting 64 agents, x8 times than before, (§$\S$5.1), limited multimodal senses (§$\S$§J) and physical needs (§$\S$§K). By incorporating these three features, our simulator fosters social dynamics (§$\S$5.2&§$\S$§O). (2) Our benchmark and datasets are challenging (§$\S$5.3&§$\S$5.4). (3) Our agents work together more effectively, with a reduced workload per agent (§$\S$5.5), and the multitasking mechanism for agents is beneficial (§$\S$5.6).
@@
In summary, our contributions are three-fold: the simulator, benchmark, and AI agent. With these contributions, we push the boundaries of multi-agent simulation by bridging the gap between virtual agents and large-scale real-world humans. This not only advances understanding of AI multi-agents but also holds potential for applications in human dynamics, social psychology, robotics, and game design. We anticipate that this work will serve as a useful foundation for the community to create new algorithms and make progress in the field of embodied AI multi-agent systems.
@@ 2 MineLand Simulator
Conventional multi-agent open-world simulators suffer from the gap between virtual agents and large-scale real-world humans. To bring this gap, we propose MineLand with three key features: large-scale scalability, limited multimodal senses, and physical needs. This section describes the design and implementation of this simulator, focusing on the architecture, observation space, state space, action space and communication. During introducing the design, we will highlight techniques that bring the three key features and omit the parts similar to other Minecraft simulators.
Table 1: Part of the comparison with related popular platforms or projects. The full table is in Table 7. Max Agents: The approximate maximum number of agents supported on a single PC (See Section 5.1 for details). Human: Whether the simulator allows humans to directly interact with AI agents. Plan as Action: Whether agents can generate plans in a specific format (e.g., code) that the simulator interprets and executes directly as actions. Sociological Experiments: Whether the simulator is designed to facilitate the study of social phenomena and emergent behavior in multi-agent systems.
Simulator Max Agents Human Plan as Action Sociological Number of Tasks
Experiments
MineLand 64+ 6000+
MineDojo [16] 1 - - - 3000+
MineRL v1.0 [3] 1 - - - 5
MarLÖ [42] 8 - - 14
Malmo [27] 8 - - -
Voyager [57] 1 - -
Smallville [40] 25 - -
2.1 Architecture
MineLand, inspired by Malmo [27], Mineflayer [43], MineDojo [16] and Voyager [57], is a Minecraft simulator where players [^2] In this work, we use the term “player” to refer to both human players and AI agents. When we mention “agent”, we specifically mean AI agents. Humans have the option to access the game either through VR or using a keyboard. can explore, and interact with each other as well as the environments. The architecture as shown in Figure 2 consists of three main modules (Details in Section C.1): the Bot Module, the Environment Module, and the Bridge Module.
What technology has caused the feature of large-scale scalability?
Existing Minecraft simulators (e.g., Malmo, MineRL, MineDojo) necessitate running a Minecraft game client for each player. This client-based approach comes with a notable drawback - it incurs substantial resource costs and most machines cannot handle running large-scale Minecraft game clients concurrently. In contrast, MineLand adopts a different approach. It simplifies each Minecraft client into a single thread, optimizing performance overhead caused by multiple clients. With MineLand, adding one more agent only requires one more thread. Based on this new technique, MineLand supports 64 or more agents on a mainstream consumer desktop PC, which is a substantial improvement compared to other Minecraft platforms that can only support up to 2 agents. We conducted relevant experiments in Section 5.1.
@@ 2.2 Observation Space
The observation space is designed to be compatible with almost all APIs of the popular MineDojo framework. MineLand provides sensor information for players: tactile information (information about the blocks surrounding the agent, which represent the objects that the agent can touch), auditory information, and visual information (RGB video from the first-person perspective of the agent). These three modalities (namely, touch, vision, and hearing) together provide the agents with multimodal senses. Note that this information is all raw perceptual information. [^3] Besides the raw perceptual information, MineLand also provides the events encountered by the agent, such as injury, death, and others. Injury events can also be regarded as tactile information, but they are presented in the form of events for simplicity.
What technology has caused the feature of limited multimodal senses?
We refer to the mechanisms of human vision and hearing, and impose limitations on the sensor perception of players, including distance attenuation, environmental obstructions, and directional constraints, to model the limited senses.
@@ 2.3 State Space
Previous simulators focus on task-oriented activities, thus the state space focuses on the inventory and equipment. We use a state space that blends task-oriented activities with the rhythms of daily life.
What technology has caused the feature of physical needs?
Basic physical needs are the foundation that leads to daily life behavior. For the rhythms of daily life, we define the states of agents for themselves: physical needs like oxygen and hunger. Blending the rhythms of daily life with task-oriented activities is what makes this simulator stand out. Imagine agents waking up in their virtual Minecraft homes, engaging in daily routines like cooking to satisfy physical food needs, but also having defined jobs (e.g., lumberjack, farmer) that involve specific task-oriented activities. This creates a natural flow between daily life and goal-driven behavior, providing a more realistic and nuanced environment for studying agent interactions and complex social dynamics.
@@ 2.4 Action Space
The simulator offers a unique action space encompassing both low-level and high-level actions. For the low-level actions, MineLand includes basic actions like walking, running, jumping, and interacting with objects. The low-level actions are the same as the traditional action space in gym-style API. We also support Reinforcement Learning based on low-level actions (For more details about experiments of Reinforcement Learning method, please refer to Section I. High-level actions, like dodging obstacles and manipulating tools, are suitable for complex tasks that consist of several low-level actions and require longer computation times. The high-level actions are implemented in the form of code, following Voyager [57]. Imagine agents navigating the world, dodging obstacles, and manipulating tools. These complex tasks generate an action sequence, allowing the simulator to continue executing the low-level actions, skip some steps earlier, or be interrupted by some special event.
Figure 3: Illustration of Tasks. We have expanded the number of programmatic tasks and creative tasks by 2 times, compared to MineDojo. Additionally, we have introduced novel hybrid tasks that combine the features of programmatic tasks and creative tasks. Customizing the number of players is supported. For multi-agents, we provide two modes: cooperative mode and competitive mode.
@@ 2.5 Communication
Diversity
Diverse communication strategies can lead to the emergence of more realistic behaviors within the simulated environment. We design three communication strategies including auditory information, body language (via visual perception), and sharing information in text media.
What communication technology has caused the feature of large-scale scalability?
For multi-agent simulators, an efficient communication mechanism is crucial, especially in situations with large-scale scalability. Traditional communication methods, like centralized broadcasting to all agents, can become computationally expensive and slow down the simulation with a large number of agents. Here we design a distance-constrained constrained communication mechanism. If an agent wants to communicate with other agents, it chats through Minecraft’s message bar. Only when the distance between other agents and the sending agent is less than a certain threshold, will other agents receive messages. Similarly, communications via auditory information and body language are limited by distance.
How is the interrupt mechanism implemented?
Most importantly, the new message is allowed to interrupt the executing code and execute this message directly before the previous code has ended. We term this as the interrupt mechanism. With this mechanism, even if an agent is working on a 5-minute extension (such as mining), it is still feasible for other agents to communicate with this working agent at any time. This interrupt mechanism was not supported in the previous work, such as Voyager [57]. Next, we will introduce how to implement it. For a high-level action, the execution of the code is divided into several steps, with each step lasting 50-200 milliseconds44450 milliseconds is the minimum time unit in Minecraft. We refer to this minimum time unit as a “tick”.. Before taking a step, the agent is provided with the running states of the previous code, either running, ready, or exceptions. After completing a step, the agent, based on the running states, can choose to either switch to a new action code or continue executing the previous code. This function of choosing is implemented by an automatic gate control system with two gates: New and Resume. New means the agent wants to switch to a new code in the following steps. Resume indicates that the agent wants to continue executing the previous code. In this way, the agent can complete a code that needs to be executed for a long period, or be interrupted at an appropriate time. You may refer Section E.4 for an example.
@@ 3 MineLand Benchmark Suite and Dataset
MineLand, as a large-scale multi-agent simulator, push the boundaries of multi-agent capabilities by enabling them to tackle complex human-like planning tasks within diverse environments. However, evaluating these advanced planning abilities necessitates sophisticated benchmarks. To address this challenge, we propose a new benchmark suit, MineLand benchmark. MineLand benchmark surpasses existing benchmarks by offering significantly more tasks (doubling programmatic and creative tasks compared to MineDojo) and introducing a novel “hybrid task” category that combines the features of programmatic and creative tasks (including Construction Tasks and Stage Performance Tasks). Additionally, this benchmark allows for flexible player numbers and exploration through cooperative and competitive modes. Competitive mode can be used to measure the differences in capabilities between different AI agents, as well as to develop adversarial learning algorithms. Refer to Section D for more details.